# Determination of the stably free cancellation property for orders

### v1

@@ 9.1. An algorithm to show that an order fails SFC
Due to the invocation of the (expensive) freeness test of [BJ11, Algorithm 3.1], Algorithm 9.1 is only practical when $\Lambda$ is an order in an algebra $A$ of ‘moderate’ dimension. However, it is also possible to obtain results when $A$ is of larger dimension as follows. A projection of $K$-algebras $A\to B$ induces a projection of $\Lambda$ onto an $\mathcal{O}$-order $\Gamma$ in $B$. If we can apply Algorithm 9.1 to show that $\Gamma$ fails SFC, then Theorem 3.1 shows that $\Lambda$ also fails SFC. Since we are free to choose $B$, we can take $B$ to be a direct sum of simple components of $A$ of small dimension, chosen such that all non-Eichler simple components of $A$ are also components of $B$ and Algorithm 9.1 finishes quickly.
@@ 9.2. Integral group rings without SFC
We now use Algorithm 9.1 and Remark 9.3 to show that certain integral group rings fail SFC. Parts of the following result have already proven by Swan [Swa83] and Chen [Che86] (see Remark 9.5).
@@
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}$, or
@@
Let $e$ be the central idempotent of $\mathbb{Q}[G]$ such that $e\mathbb{Q}[G]$ is the direct sum of all simple components of $\mathbb{Q}[G]$ that are either commutative or totally definite quaternion algebras. Then $e\mathbb{Z}[G]$ fails SFC. Moreover, $\mathbb{Z}[G]$ fails SFC but $\mathbb{Z}[G/N]$ has SFC for every proper quotient $G/N$.
@@
For each group $G$ listed, verification that $e\mathbb{Z}[G]$ fails SFC was performed using Algorithm 9.1. Theorem 3.1 then implies that $\mathbb{Z}[G]$ also fails SFC.
@@
then $\mathbb{Z}[H]$ has LFC by Theorem 4.2. Therefore it suffices to consider maximal quotients $H$ such that $\mathbb{Q}[H]$ does not satisfy the Eichler condition and is not isomorphic to any of the groups in (9.1). The only groups $G$ listed that have such maximal quotients are $\tilde{T}\times C_{2}^{2}$, $G_{(32,14)}$ and $G_{(64,14)}$. For $G=\tilde{T}\times C_{2}^{2}$, each such $H$ is isomorphic to $\tilde{T}\times C_{2}$, and so $\mathbb{Z}[H]$ has SFC by Theorem 6.6. For $G=G_{(32,14)}$, each such $H$ is isomorphic to $G_{(16,4)}\cong C_{4}\rtimes C_{4}$ and so $\mathbb{Z}[H]$ has SFC by Theorem 4.11. For $G=G_{(64,14)}$, each such $H$ is isomorphic to $G_{(32,10)}\cong Q_{8}\rtimes C_{4}$. Since $G_{(32,10)}$ has a quotient isomorphic to $Q_{16}$ and $m_{\mathbb{H}}(G_{(32,10)})=m_{\mathbb{H}}(Q_{16})=2$, the group ring $\mathbb{Z}[G_{(32,10)}]$ has SFC by Theorem 4.10. ∎
@@
Let $G$ be a group listed in Theorem 9.4. Then Corollary 4.4 implies that every maximal $\mathbb{Z}$-order $\mathcal{M}$ in $\mathbb{Q}[G]$ has LFC. Thus if such an $\mathcal{M}$ contains $\mathbb{Z}[G]$, then $e\mathcal{M}$ is a maximal $\mathbb{Z}$-order containing $e\mathbb{Z}[G]$, and by Theorem 3.1 $e\mathcal{M}$ also has LFC (and thus SFC). Moreover, $e\mathbb{Z}[G]$ is not maximal. Therefore it is not necessary to perform Step (2) of Algorithm 9.1 when applied to $e\mathbb{Z}[G]$.
@@
This follows from Lemma 4.1, Theorem 4.2 and Theorem 9.4. ∎
Remark 9.8.
The following question naturally arises: for a finite group $G$ and $e$ as in the statement of Theorem 9.4, is it always true that $\mathbb{Z}[G]$ has SFC if and only if $e\mathbb{Z}[G]$ has SFC? One direction is clear by Theorem 3.1. If the other direction could be proven, then this would make the problem of determining whether $\mathbb{Z}[G]$ has SFC much easier from an algorithmic point of view. Of course, one can also ask the analogous question for LFC.
Remark 9.9.
Let $G$ be a finite group and let $\tilde{e}$ be the central idempotent of $\mathbb{Q}[G]$ such that $\tilde{e}\mathbb{Q}[G]$ is the direct sum of all simple components of $\mathbb{Q}[G]$ that are totally definite quaternion algebras. Then for $G=Q_{16}\times C_{2}$, the order $\tilde{e}\mathbb{Z}[G]$ fails SFC, but for $G=Q_{12}\times C_{2}$, it does have SFC. Thus $e$ cannot be replaced by $\tilde{e}$ for all groups listed in Theorem 9.4.
@@ 10.1. Setup
In particular, this is a locally free $\Lambda$-lattice of rank $1$. We refer to such a lattice $M(u)$ as a Reiner–Ullom module. Recall from Corollary 5.3 that we have a group homomorphism
@@
We shall relate Reiner–Ullom modules to generalised Swan modules and use this to give an alternative description of $\ker(\partial)$. Under the further assumption that $A_{2}$ is Eichler, this will lead to an algorithm to test whether $\Lambda$ has SFC or not.
10.2. Generalised Swan modules and Reiner–Ullom modules
For $i=1,2$, let $\mathcal{M}_{i}$ be a choice of $\mathcal{O}$-order in $A_{i}$ containing $\Lambda_{i}$ (such choices exists by [Rei03, (10.4)]) and let $\mathfrak{f}_{i}$ be any proper full two-sided ideal of $\mathcal{M}_{i}$ that is contained in $\Lambda\cap A_{i}$. Let $\mathcal{M}=\mathcal{M}_{1}\oplus\mathcal{M}_{2}$ and let $\mathfrak{f}=\mathfrak{f}_{1}\oplus\mathfrak{f}_{2}$. For $i=1,2$, let $C_{i}$ be the centre of $A_{i}$ and let $\mathfrak{g}_{i}:=\mathfrak{f}_{i}\cap\mathcal{O}_{C_{i}}=\mathfrak{f}_{i}\cap C_{i}$. Let $C=C_{1}\oplus C_{2}$ and let $\mathfrak{g}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}$. We set
@@ 10.2. Generalised Swan modules and Reiner–Ullom modules
Since for any $x_{1}\in\Lambda_{1}$ with $(x_{1},0)\in\Lambda$ we have $x_{1}\in\ker(g_{1})$, it follows that $\mathfrak{f}_{1}\subseteq\mathfrak{h}_{1}$. Similarly, $\mathfrak{f}_{2}\subseteq\mathfrak{h}_{2}$ and therefore $\mathfrak{f}\subseteq\mathfrak{h}$. Recall from Definition 8.1 that for $\beta\in\mathcal{M}\cap A^{\times}$, the generalised Swan module $X_{\beta}=X_{\beta,\Lambda,\mathcal{M}}$ is defined to be $\mathcal{M}\beta\cap\Lambda$.
The following result shows that under mild assumptions, Reiner–Ullom modules are in fact (isomorphic to) generalised Swan modules.
@@ 10.3. Alternative description of $\ker(\partial)$
Let $\iota_{2}\colon(\Lambda_{2}/\mathfrak{f}_{2})^{\times}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$ and $\theta_{2}\colon(\Lambda_{2}/\mathfrak{f}_{2})^{\times}\longrightarrow(\Lambda_{2}/\mathfrak{h}_{2})^{\times}$ be the canonical group homomorphisms (recall from §10.2 that $\mathfrak{f}_{2}\subseteq\mathfrak{h}_{2}$). Then $\iota_{2}$ is clearly injective and $\theta_{2}$ is surjective by [Bas68, Chapter III, (2.9) Corollary]. Moreover, it follows from the definition of $\nu_{2}^{\prime}$ that $\tilde{\mu}\circ\tilde{\nu}\circ\iota_{2}=\theta_{2}\circ\nu_{2}^{\prime}$. Every element in $\ker(\theta_{2})$ is represented by an element of $\Lambda_{2}$ of the form $\lambda_{2}=1+h_{2}$ with $h_{2}\in\mathfrak{h}_{2}$. Since
@@
Let $x_{2}\in\left(\Lambda_{2}/\mathfrak{h}_{2}\right)^{\times}$ and let $\alpha_{2}$ be as in the statement of the lemma. Then we have
@@ 10.4. An algorithm for determining SFC using fibre products
    Compute $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$ and $\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})$ as finite abelian groups, where $\mathfrak{g}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}$.
@@
    Compute elements $\overline{\gamma}_{1},\ldots,\overline{\gamma}_{r}\in\overline{\Lambda}_{2}^{\times}$ such that their images are generators of the subgroup $\ker(\nu_{2}^{\prime})/[\overline{\Lambda}_{2}^{\times},\overline{\Lambda}_{2}^{\times}]\leq\overline{\Lambda}_{2}^{\times\mathrm{ab}}$.
@@
By Lemma 10.2, we have $\ker(\partial)={\tilde{g}_{2}}(\ker(\nu^{\prime}_{2}))$, where $\tilde{g}_{2}:\Lambda_{2}/\mathfrak{h}_{2}\cong\overline{\Lambda}$ is the ring isomorphism induced by $g_{2}$. Thus $\hat{g}_{2}^{-1}(\overline{\gamma}_{1}),\ldots,\hat{g}_{2}^{-1}(\overline{\gamma}_{r})$ is a set of generators of $\ker(\partial)/[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$. Moreover, by Proposition 10.1, for each $i\in\{1,\ldots,r\}$, we have $X_{(1,\beta_{i})}\cong M(g_{2}(\beta_{i})^{-1})$ as $\Lambda$-lattices. Therefore the correctness of the output in step (16) follows from Corollary 5.12 with $N=[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$. ∎
@@
Step (8) is straightforward and Step (9) can be performed using the methods outlined in §11.3. Note that, in practice, Step (9) is often the bottleneck in the computation. Steps (10), (13), and (14) can be performed using standard methods in computational group theory, as described in [HEO05], for example, and the explicit definition of $\nu_{2}^{\prime}$ given in (10.3). Step (15) is straightforward and in Step (16), the freeness testing can be performed either as described in Step (11) of Algorithm 8.9 or the (more efficient) method described in §12. ∎
@@ 11.2. Primary decomposition and generators of unit groups of quotient rings
(i) As in the proof of [Neu99, (3.4) Lemma] one can show that there is a product of maximal ideals $\mathfrak{q}_{i}$ of $\Gamma$ such that $\mathfrak{q}_{1}\cdots\mathfrak{q}_{m}\subseteq\mathfrak{g}$. Since $\mathfrak{q}_{1}\cdots\mathfrak{q}_{m}\subseteq\mathfrak{g}\subseteq\mathfrak{p}_{i}$ for $i=1,\ldots,r$ each of the maximal ideals $\mathfrak{q}_{i}$ is one of $\mathfrak{p}_{1},\ldots,\mathfrak{p}_{r}$. Hence there exist $e_{1},\ldots,e_{r}\in\mathbb{Z}_{>0}$ such that $\mathfrak{p}_{1}^{e_{1}}\cdots\mathfrak{p}_{r}^{e_{r}}\subseteq\mathfrak{g}$.
(ii) Since $\mathfrak{p}_{i}^{e_{i}}\subseteq\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}\subseteq\mathfrak{p}_{i}$, the radical of $\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}$ is the maximal ideal $\mathfrak{p}_{i}$. This implies that $\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}$ is $\mathfrak{p}_{i}$-primary ([AM69, Prop. (4.2)]).
(iii) It suffices to show that $\mathfrak{g}=\bigcap_{i=1}^{r}(\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g})=\prod_{i=1}^{r}(\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g})$. The second equality is clear because the ideals $\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}$ are pairwise coprime. For the first equality the inclusion $\subseteq$” is obvious. Conversely, $\prod_{i=1}^{r}(\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g})\subseteq\mathfrak{p}_{1}^{e_{1}}\cdots\mathfrak{p}_{r}^{e_{r}}+\mathfrak{g}=\mathfrak{g}$. ∎
@@
Let $\mathcal{O}_{C}$ be the (unique) maximal $\mathcal{O}$-order in $C$. This is equal to the integral closure of $\Gamma$ in $C$. By the going-up theorem each of the maximal ideals $\mathfrak{p}_{i}$ is of the form $\mathfrak{P}_{i}\cap\Gamma$ for a maximal ideal $\mathfrak{P}_{i}$ of $\mathcal{O}_{C}$. Let $\{\mathfrak{P}_{1},\ldots,\mathfrak{P}_{n}\}$ be the set of maximal ideals of $\mathcal{O}_{C}$ with $\mathfrak{g}\mathcal{O}_{C}\subseteq\mathfrak{P}_{i}$. Then
@@
Let $C$ be be an étale $K$-algebra and let $\Gamma$ be an $\mathcal{O}$-order in $C$. Let $\mathfrak{g}$ be non-trivial proper ideal of $\Gamma$ of full rank. Then there exists an algorithm that computes the primary decomposition of $\mathfrak{g}$.

### v2

@@ 9.1. An algorithm to show that an order fails SFC
Due to the invocation of the (expensive) freeness test of [BJ11, Algorithm 3.1], Algorithm 9.1 is only practical when $\Lambda$ is an order in an algebra $A$ of ‘moderate’ dimension. However, it is also possible to obtain results when $A$ is of larger dimension as follows. A projection of $K$-algebras $A\to B$ induces a projection of $\Lambda$ onto an $\mathcal{O}$-order $\Gamma$ in $B$. If we can apply Algorithm 9.1 to show that $\Gamma$ fails SFC, then Theorem 3.1 shows that $\Lambda$ also fails SFC. Since we are free to choose $B$, we can take $B$ to be a direct sum of simple components of $A$ of small dimension, chosen such that all simple components of $A$ that do not satisfy the Eichler condition are also components of $B$ and Algorithm 9.1 finishes quickly.
@@ 9.2. Integral group rings without SFC
We now use Algorithm 9.1 and Remark 9.3 to show that certain integral group rings fail SFC. Parts of the following result, which is Theorem 1.1 from the introduction, have already been proven by Swan [Swa83] and Chen [Che86] (see Remark 9.5).
@@
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}^{2}$, or
@@
Remark 9.6.
In order, the groups listed in Theorem 9.4 (i) are isomorphic to $G_{(16,12)}$, $G_{(24,12)}$, $G_{(32,41)}$, $G_{(40,7)}$, and those in (ii) to $G_{(96,198)}$, $G_{(96,188)}$, $G_{(480,960)}$.
@@
We first show that $\mathbb{Z}[G]$ fails SFC. Let $e$ be the central idempotent of $\mathbb{Q}[G]$ such that $e\mathbb{Q}[G]$ is the direct sum of all simple components of $\mathbb{Q}[G]$ that are either commutative or totally definite quaternion algebras. For each group $G$ listed apart from $G=\tilde{O}\times C_{2}$, verification that $e\mathbb{Z}[G]$ fails SFC was performed using Algorithm 9.1. Theorem 3.1 then implies that $\mathbb{Z}[G]$ also fails SFC. For $G=\tilde{O}\times C_{2}$ there exists a normal subgroup $N$ of order $2$ such that $G/N\cong S_{4}\times C_{2}$. For this choice of $G$ and $N$, the verification that the order $\Gamma:=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$ fails SFC was performed using Algorithm 9.1; again, Theorem 3.1 then implies that $\mathbb{Z}[G]$ also fails SFC.
@@
then $\mathbb{Z}[H]$ has LFC by Theorem 4.2. Therefore it suffices to consider maximal quotients $H$ such that $\mathbb{Q}[H]$ does not satisfy the Eichler condition and is not isomorphic to any of the groups in (9.1). The only groups $G$ listed that have such maximal quotients are $\tilde{T}\times C_{2}^{2}$, $\tilde{I}\times C_{2}^{2}$, $G_{(32,14)}$ and $G_{(64,14)}$. For $G=\tilde{T}\times C_{2}^{2}$, each such $H$ is isomorphic to $\tilde{T}\times C_{2}$, and so $\mathbb{Z}[H]$ has SFC by Theorem 6.6. For $G=\tilde{I}\times C_{2}^{2}$, each such $H$ is isomorphic to $\tilde{I}\times C_{2}$, and so $\mathbb{Z}[H]$ has SFC by Theorem 13.1. For $G=G_{(32,14)}$, each such $H$ is isomorphic to $G_{(16,4)}\cong C_{4}\rtimes C_{4}$ and so $\mathbb{Z}[H]$ has SFC by Theorem 4.13. For $G=G_{(64,14)}$, each such $H$ is isomorphic to $G_{(32,10)}\cong Q_{8}\rtimes C_{4}$. Since $G_{(32,10)}$ has a quotient isomorphic to $Q_{16}$ and $m_{\mathbb{H}}(G_{(32,10)})=m_{\mathbb{H}}(Q_{16})=2$, the group ring $\mathbb{Z}[G_{(32,10)}]$ has SFC by Theorem 4.12. ∎
@@
Let $G$ be any group listed in Theorem 9.4 and let $\mathcal{M}$ be any maximal $\mathbb{Z}$-order in $\mathbb{Q}[G]$. Then every binary polyhedral quotient of $G$ is isomorphic to $Q_{8}$, $Q_{12}$, $Q_{16}$, $Q_{20}$, $\tilde{T}$, $\tilde{O}$, or $\tilde{I}$, and hence $\mathcal{M}$ has LFC (and thus SFC) by Corollary 4.6.
@@
This follows from Theorem 9.4 and Corollary 4.3. ∎
@@ 10.1. Setup
In particular, this is a locally free $\Lambda$-lattice of rank $1$. We refer to such a lattice $M(u)$ as a pullback lattice. Recall from Corollary 5.3 that we have a group homomorphism
@@
We shall relate pullback lattices to test lattices and use this to give an alternative description of $\ker(\partial)$. Under the further assumption that $A_{1}$ satisfies the Eichler condition, this will lead to an algorithm to test whether $\Lambda$ has SFC or not.
10.2. Test lattices and pullback lattices
For $i=1,2$, let $\mathcal{M}_{i}$ be a choice of $\mathcal{O}$-order in $A_{i}$ containing $\Lambda_{i}$ (such choices exist by [Rei03, (10.4)]) and let $\mathfrak{f}_{i}$ be any proper full two-sided ideal of $\mathcal{M}_{i}$ that is contained in $\Lambda\cap A_{i}$. Let $\mathcal{M}=\mathcal{M}_{1}\oplus\mathcal{M}_{2}$ and let $\mathfrak{f}=\mathfrak{f}_{1}\oplus\mathfrak{f}_{2}$. For $i=1,2$, let $C_{i}$ be the centre of $A_{i}$ and let $\mathfrak{g}_{i}:=\mathfrak{f}_{i}\cap\mathcal{O}_{C_{i}}=\mathfrak{f}_{i}\cap C_{i}$. Let $C=C_{1}\oplus C_{2}$ and let $\mathfrak{g}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}$. We set
@@ 10.2. Test lattices and pullback lattices
Since for any $x_{1}\in\Lambda_{1}$ with $(x_{1},0)\in\Lambda$ we have $x_{1}\in\ker(g_{1})$, it follows that $\mathfrak{f}_{1}\subseteq\mathfrak{h}_{1}$. Similarly, $\mathfrak{f}_{2}\subseteq\mathfrak{h}_{2}$ and therefore $\mathfrak{f}\subseteq\mathfrak{h}$. Recall from Definition 8.1 that for $\beta\in\mathcal{M}\cap A^{\times}$, the test lattice $X_{\beta}=X_{\beta,\Lambda,\mathcal{M}}$ is defined to be $\mathcal{M}\beta\cap\Lambda$.
The following result shows that under mild assumptions, pullback lattices are in fact (isomorphic to) test lattices.
@@ 10.3. Alternative description of $\ker(\partial)$
Let $\iota_{2}\colon(\Lambda_{2}/\mathfrak{f}_{2})^{\times}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$ and $\theta_{2}\colon(\Lambda_{2}/\mathfrak{f}_{2})^{\times}\longrightarrow(\Lambda_{2}/\mathfrak{h}_{2})^{\times}$ be the canonical group homomorphisms (recall from §10.2 that $\mathfrak{f}_{2}\subseteq\mathfrak{h}_{2}$). Then $\iota_{2}$ is clearly injective and $\theta_{2}$ is surjective by [Bas68, Chapter III, (2.9) Corollary]. Moreover, it follows from the definition of $\nu_{2}^{\prime}$ that $\tilde{\mu}\circ\tilde{\nu}\circ\iota_{2}=\nu_{2}^{\prime}\circ\theta_{2}$. Every element in $\ker(\theta_{2})$ is represented by an element of $\Lambda_{2}$ of the form $\lambda_{2}=1+h_{2}$ with $h_{2}\in\mathfrak{h}_{2}$. Since
@@
Let $x_{2}\in\left(\Lambda_{2}/\mathfrak{h}_{2}\right)^{\times}$ and let $\alpha_{2}$ be as in the definition of the map $\nu_{2}^{\prime}$. Then we have
@@ 10.4. An algorithm for determining SFC using fibre products
    Compute $I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$ and $\tilde{\nu}((\Lambda/\mathfrak{f})^{\times})\leq I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}$ as finite abelian groups, where $\mathfrak{g}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}$.
@@
    Compute elements $\overline{\gamma}_{1},\ldots,\overline{\gamma}_{r}\in\overline{\Lambda}_{2}^{\times}$ such that the set of their images in $\overline{\Lambda}_{2}^{\times\mathrm{ab}}$ is a minimal generating set of the subgroup $\ker(\nu_{2}^{\prime})/[\overline{\Lambda}_{2}^{\times},\overline{\Lambda}_{2}^{\times}]\leq\overline{\Lambda}_{2}^{\times\mathrm{ab}}$.
@@
By Lemma 10.2, we have $\ker(\partial)={\tilde{g}_{2}}(\ker(\nu^{\prime}_{2}))$, where $\tilde{g}_{2}:\Lambda_{2}/\mathfrak{h}_{2}\cong\overline{\Lambda}$ is the ring isomorphism induced by $g_{2}$. Thus the set of images of $\tilde{g}_{2}(\overline{\gamma}_{1}),\ldots,\tilde{g}_{2}(\overline{\gamma}_{r})$ in $\overline{\Lambda}^{\times\mathrm{ab}}=\overline{\Lambda}^{\times}{}{/}{}[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$ generates the subgroup $\ker(\partial)/[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]\leq\overline{\Lambda}^{\times\mathrm{ab}}$. Moreover, by Proposition 10.1, for each $i\in\{1,\ldots,r\}$, we have $X_{(1,\beta_{i})}\cong M(g_{2}(\beta_{i})^{-1})$ as $\Lambda$-lattices. Therefore the correctness of the output in step (16) follows from Corollary 5.12 with $N=[\overline{\Lambda}^{\times},\overline{\Lambda}^{\times}]$. ∎
@@
Step (8) is straightforward and Step (9) can be performed using the methods outlined in §11.3. Note that, in practice, Step (9) is often the bottleneck in the computation. Steps (10), (13), and (14) can be performed using standard methods in computational group theory, as described in [HEO05], for example, and the explicit definition of $\nu_{2}^{\prime}$ given in (10.3). Step (15) is straightforward and in Step (16), the freeness testing can be performed either as described in Step (11) of Algorithm 8.9 or the method described in §12. ∎
@@ 11.2. Primary decomposition and generators of unit groups of quotient rings
(i) As in the proof of [Neu99, (3.4) Lemma] one can show that there is a product of maximal ideals $\mathfrak{q}_{i}$ of $\Gamma$ such that $\mathfrak{q}_{1}\cdots\mathfrak{q}_{m}\subseteq\mathfrak{g}$. We take $m$ to be minimal. Suppose that one of the $\mathfrak{q}_{i}$, say $\mathfrak{q}_{1}$, is not contained in the set $\{\mathfrak{p}_{1},\ldots,\mathfrak{p}_{r}\}$. Then we claim that $\mathfrak{q}_{2}\cdots\mathfrak{q}_{m}\subseteq\mathfrak{g}$ contradicting the minimality of $m$. Indeed, we then have $\mathfrak{q}_{1}+\mathfrak{g}=\Gamma$, so that we can write $1=x+y$ with $x\in\mathfrak{q}_{1}$ and $y\in\mathfrak{g}$. For any $z\in\mathfrak{q}_{2}\cdots\mathfrak{q}_{m}$ we obtain $z=zx+zy\in\mathfrak{g}$. We conclude that there exist $e_{1},\ldots,e_{r}\in\mathbb{Z}_{>0}$ such that $\mathfrak{p}_{1}^{e_{1}}\cdots\mathfrak{p}_{r}^{e_{r}}\subseteq\mathfrak{g}$.
(ii) Since $\mathfrak{p}_{i}^{e_{i}}\subseteq\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}\subseteq\mathfrak{p}_{i}$, the radical of $\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}$ is the maximal ideal $\mathfrak{p}_{i}$. This implies that $\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}$ is $\mathfrak{p}_{i}$-primary by [AM69, Proposition (4.2)].
(iii) By part (ii), it suffices to show that $\mathfrak{g}=\bigcap_{i=1}^{r}(\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g})=\prod_{i=1}^{r}(\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g})$. The second equality is clear because the ideals $\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g}$ are pairwise coprime. For the first equality, observe that $\prod_{i=1}^{r}(\mathfrak{p}_{i}^{e_{i}}+\mathfrak{g})\subseteq\mathfrak{p}_{1}^{e_{1}}\cdots\mathfrak{p}_{r}^{e_{r}}+\mathfrak{g}=\mathfrak{g}$; the reverse inclusion is obvious. ∎
@@
Let $\mathcal{O}_{C}$ be the (unique) maximal $\mathcal{O}$-order in $C$. This is equal to the integral closure of $\Gamma$ in $C$. By the going-up theorem each of the maximal ideals $\mathfrak{p}_{i}$ is of the form $\mathfrak{P}_{i}\cap\Gamma$ for some maximal ideal $\mathfrak{P}_{i}$ of $\mathcal{O}_{C}$. Let $\{\mathfrak{P}_{1},\ldots,\mathfrak{P}_{n}\}$ be the set of maximal ideals of $\mathcal{O}_{C}$ with $\mathfrak{g}\mathcal{O}_{C}\subseteq\mathfrak{P}_{i}$. Then
@@
Let $C$ be an étale $K$-algebra and let $\Gamma$ be an $\mathcal{O}$-order in $C$. Let $\mathfrak{g}$ be a proper ideal of $\Gamma$ of full rank. Then there exists an algorithm that computes the primary decomposition of $\mathfrak{g}$.